be and ? = ordered bases for R Let T be a linear transformation such that x1 -x 2x1 + x2 X1 x2 Then ITPob What is a b? Answer: Check
Answer #1

Solution:

\small T\begin{bmatrix} 2\\ -1\\ \end{bmatrix}=\begin{bmatrix} 3\\ 3\\ \end{bmatrix}= 3\begin{bmatrix} 1\\ 1\\ \end{bmatrix}+0\begin{bmatrix} 2\\ -3\\ \end{bmatrix}

\small T\begin{bmatrix} 1\\ 2\\ \end{bmatrix}=\begin{bmatrix} -1\\ 4\\ \end{bmatrix}= 1\begin{bmatrix} 1\\ 1\\ \end{bmatrix}+\left ( -1 \right )\begin{bmatrix} 2\\ -3\\ \end{bmatrix}

\small \therefore \left [ T \right ]^{\Omega }_{\Gamma }=\begin{bmatrix} 3 &1 \\ 0 & -1 \end{bmatrix}

Comparing with \small \left [ T \right ]^{\Omega }_{\Gamma }=\begin{bmatrix} a &1 \\ 0 & b \end{bmatrix} we have

\small a=3\: \: \: ,\: \: \: b=-1

\small {\color{DarkBlue} \therefore a+b=3+\left ( -1 \right )=2}